# KramaBench: A Benchmark for AI Systems on Data Intensive Tasks

### v2

@@ Appendix A Extended Experiment Results
| Archaeology | Astronomy | Biomedical | Environment | Legal | Wildfire | Total | ||
| DS-Guru no-context | GPT-o3 | 17.83% | 12.93% | 19.48% | 19.17% | 9.94% | 16.13% | 14.93% |
| GPT-4o | 15.09% | 9.15% | 12.16% | 11.26% | 8.88% | 7.15% | 10.05% | |
| Claude-3.5 | 16.52% | 10.63% | 9.87% | 12.51% | 9.80% | 0.00% | 11.63% | |
| Llama3-3Instruct | 14.44% | 12.17% | 10.24% | 10.35% | 8.20% | 8.06% | 9.93% | |
| DeepSeek-R1 | 18.79% | 8.53% | 8.25% | 12.71% | 11.39% | 8.90% | 11.56% | |
| Qwen2-5Coder | 10.24% | 6.74% | 7.71% | 7.14% | 1.52% | 4.53% | 6.62% | |
| DS-Guru one-shot | GPT-o3 | 23.90% | 21.14% | 18.29% | 28.48% | 18.49% | 25.08% | 22.85% |
| GPT-4o | 14.26% | 10.58% | 9.38% | 20.37% | 10.96% | 19.21 | 14.86% | |
| Claude-3-5 | 17.07% | 10.24% | 9.44% | 22.27% | 11.47% | 17.93% | 15.48% | |
| Llama3-3Instruct | 8.92% | 10.44% | 4.45% | 12.44% | 8.64% | 12.90% | 10.23% | |
| DeepSeek-R1 | 16.78% | 15.23% | 8.06% | 14.23% | 11.89% | 9.65% | 12.64% | |
| Qwen2-5Coder | 9.72% | 11.57% | 5.37% | 15.13% | 8.96% | 13.22% | 11.26% | |
| DS-Guru few-shot | GPT-o3 | 27.78% | 23.22% | 19.56% | 33.67% | 35.14% | 32.53% | 31.92% |
| GPT-4o | 18.97% | 19.29% | 12.51% | 27.14% | 25.23% | 26.07% | 23.60% | |
| Claude-3-5 | 16.24% | 14.02% | 14.80% | 33.83% | 26.36% | 25.02% | 24.22% | |
| Llama3-3Instruct | 15.57% | 13.85% | 11.63% | 19.37% | 15.57% | 21.56% | 17.11% | |
| DeepSeek-R1 | 22.29% | 10.79% | 9.65% | 15.45% | 11.75% | 10.76% | 13.37% | |
| Qwen2-5Coder | 11.83% | 14.91% | 7.51% | 18.39% | 13.70% | 18.51% | 15.15% | |
| smolagents DR | GPT-o3 | 41.67% | 25% | 44.44% | 45% | 44.83% | 47.62% | 44.45% |
| GPT-4o | 25% | 25% | 22.22% | 20% | 56.67% | 38.1% | 39% | |
| Claude-3-5 | 16.67% | 25% | 33.33% | 25% | 66.66% | 66.66% | 47% | |
| Claude-3-7 | 41.67% | 33.33% | 77.78% | 80% | 63.33% | 71.43% | 59% | |
@@
| System | Metric | Archaeology | Astronomy | Biomedical | Environment | Legal | Wildfire | Total |
| DS-Guru few-shot | Score | 25.00% | 3.17% | 2.71% | 17.02% | 16.25% | 49.42% | 21.78% |
| (GPT-o3) | Avg. runtime/task (min) | 0.47 | 0.49 | 0.43 | 0.83 | 1.44 | 0.81 | 0.76 |
| smolagents DR | Claude-3-7 | 33.33% | 33.33% | 44.44% | 65% | 63.33% | 66.67% | 57.85% |
| Avg. runtime/task (min) | 2.22 | 5.13 | 40.38 | 3.72 | 2.12 | 2.11 | 6.10 | |
| OpenAI DR | Score | 40% | 33.33% | 44.45% | 61.67% | 50% | 67.28% | 52.18% |
@@ B.2 Ablation Studies
| SUT | Archaeology | Astronomy | Biomedical | Environment | Legal | Wildfire | Overall | Runtime |
@@
| SUT | Archaeology | Astronomy | Biomedical | Environment | Legal | Wildfire | Overall | Runtime |
@@
Appendix C Dataset Details
@@
Archaeology: the data files consists of chronological, archaeological, faunal, and botanical data supporting the presence of Holocene hunter-gatherers on the Maltese Islands in the Mediterranean from roughly 8000 years ago to 7500 years ago. The files were collected from the publicly available data associated with the papers Groucutt et al. (2021); Scerri et al. (2025).
Astronomy: the data files consist of the OMNI dataset Papitashvili & King (2020a, b) that contains near-Earth solar wind, plasma, and magnetic field data, the Swarm dataset Siemes et al. (2016); European Space Agency (2013) that contains the magnetic field and geomagnetic field data, the SILSO Sunspot Number data Clette & Lefèvre (2015), Space-Track.org Two-Line Element Sets (TLEs) U.S. Space Command (2025), the National Oceanic and Atmospheric Administration (NOAA) Flux Forecast dataset U.S. Air Force & NOAA Space Weather Prediction Center (2025), and NOAA GOES Satellite dataset NOAA Office of Satellite and Product Operations (1994). The combination of these datasets has been used to analyze how activity from the Sun affects Earth’s atmosphere, ocean currents, and weather by the authors of Briden et al. (2023); Parker & Linares (2024).
@@
Environment: the data files consist of beach water quality dataset from Massachusetts Environment Public Health Tracking (EPHT) Massachusetts Department of Public Health (2025b), the Massachusetts Bay beach dataset from Massachusetts Water Resources Authority (MWRA) Massachusetts Water Resources Authority (2025b), and the rainfall dataset from NOAA National Weather Service National Weather Service (2025), from 2002 to 2025. The data has been used in yearly reports Massachusetts Department of Public Health (2025a); Massachusetts Water Resources Authority (2025a) to uncover trends in beach water pollution and the correlation between rainfall and water quality.
@@
Legal: the datasets consists of 136 data files, accessible through the Federal Trade Commission (FTC) portal Federal Trade Commission (2025b) and Wikipedia Wikipedia contributors (2025), including information on merger filings, civil penalty actions, etc. The data is used in visualizations and dashboards that analyze nation-level debt collection and fraud detection, available at Federal Trade Commission (2025c, a).
@@
Wildfire: the datasets consists of NOAA wildfire dataset National Centers for Environmental Information (2025) (NCEI), National Interagency Fire Center (NIFC) Fire Information National Interagency Fire Center (2025), US Environmental Protection Agency (EPA) Air Quality Annual Data U.S. Environmental Protection Agency (2025), US Election 2020 Dataset Fontes (2020), Zillow Home Value Index Dataset Robikscube (2021), US Census 2020 U.S. Census Bureau (2025), and the Large wildfire Incident Status Summary Young et al. (2021) to understand wildfire incident location, cause, and consequences in the US from 2002 to 2016. This data has been used for analysis in the reports published by the NOAA and NIFC NCEI.Monitoring.Info@noaa.gov (2025); Center .
Appendix D Task Details
@@
| Archeology | 12 | 71 | 50.00% | 5 | 2 | 7.5MB |
| Astronomy | 12 | 68 | 50.00% | 1556 | 8 | 486MB |
| Biomedical | 9 | 38 | 66.66% | 7 | 2 | 175MB |
| Environment | 20 | 148 | 70.00% | 37 | 3 | 31MB |
| Legal | 30 | 188 | 53.33% | 136 | 2 | 1.3MB |
| Wildfire | 21 | 120 | 71.42% | 23 | 7 | 1GB |
| Total | 104 | 633 | 60.58% | 1764 | 24 | 1.7GB |
@@
Appendix E Evaluation Details

### v3

@@ Appendix A Extended Experiment Results
| Archeology | Astronomy | Biomedical | Environment | Legal | Wildfire | Total | ||
| DS-Guru no-context | GPT-o3 | 11.11 ± 3.93 | 10.00 ± 8.16 | 11.11 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 10.17 ± 0.92 | 5.36 ± 1.20 |
| GPT-4o | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | |
| Llama-3.3-Instruct | 5.56 ± 3.93 | 0.00 ± 0.00 | 11.11 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 1.59 ± 2.24 | 1.96 ± 0.80 | |
| Deepseek-R1 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | |
| Qwen2-5Coder* | 10.24 | 6.74 | 7.71 | 7.14 | 1.52 | 4.53 | 6.62 | |
| Claude-3.5* | 16.52 | 10.63 | 9.87 | 12.51 | 9.80 | 0.00 | 11.63 | |
| Claude-3.7 | 4.17 ± 4.17 | 10.00 ± 0.00 | 0.00 ± 0.00 | 1.58 ± 1.58 | 1.67 ± 1.67 | 8.96 ± 0.57 | 4.11 ± 0.79 | |
| DS-Guru one-shot | GPT-o3 | 8.33 ± 6.80 | 16.67 ± 4.71 | 7.51 ± 5.31 | 23.89 ± 6.71 | 15.56 ± 1.57 | 42.80 ± 3.48 | 21.35 ± 0.84 |
| GPT-4o | 11.11 ± 7.86 | 13.33 ± 4.71 | 5.19 ± 7.33 | 12.33 ± 3.30 | 6.67 ± 0.00 | 19.85 ± 3.61 | 11.54 ± 1.47 | |
| Llama-3.3-Instruct | 0.00 ± 0.00 | 6.67 ± 4.71 | 7.41 ± 5.24 | 0.00 ± 0.00 | 0.00 ± 0.00 | 26.37 ± 5.02 | 6.74 ± 1.49 | |
| Deepseek-R1 | 2.78 ± 3.93 | 6.67 ± 4.71 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.98 ± 0.80 | |
| Qwen2-5Coder* | 9.72 | 11.57 | 5.37 | 15.13 | 8.96 | 13.22 | 11.26 | |
| Claude-3.5* | 17.07 | 10.24 | 9.44 | 22.27 | 11.47 | 17.93 | 15.48 | |
| Claude-3.7 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 6.67 ± 1.67 | 3.33 ± 3.33 | 4.76 ± 0.00 | 3.27 ± 1.31 | |
| DS-Guru few-shot | GPT-o3 | 16.67 ± 0.00 | 26.67 ± 4.71 | 7.41 ± 5.24 | 50.11 ± 4.09 | 61.03 ± 5.78 | 52.02 ± 3.79 | 43.71 ± 1.94 |
| GPT-4o | 13.89 ± 3.93 | 20.00 ± 0.00 | 3.70 ± 5.24 | 18.44 ± 6.00 | 37.78 ± 4.16 | 36.67 ± 3.53 | 26.20 ± 2.34 | |
| Llama-3.3-Instruct | 13.89 ± 3.93 | 10.00 ± 0.00 | 11.11 ± 0.00 | 0.00 ± 0.00 | 6.67 ± 0.00 | 29.72 ± 1.27 | 11.67 ± 0.65 | |
| Deepseek-R1 | 5.56 ± 3.93 | 6.67 ± 4.71 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 1.31 ± 0.46 | |
| Qwen2-5Coder* | 11.83 | 14.91 | 7.51 | 18.39 | 13.70 | 18.51 | 15.15 | |
| Claude-3.5* | 16.24 | 14.02 | 14.80 | 33.83 | 26.36 | 25.02 | 24.22 | |
| Claude-3.7 | 20.83 ± 4.17 | 25.00 ± 15.00 | 0.00 ± 0.00 | 9.17 ± 9.17 | 31.67 ± 1.67 | 18.17 ± 5.20 | 19.75 ± 3.36 | |
| smolagents DR | GPT-o3 | 27.78 ± 7.86 | 26.67 ± 4.71 | 37.04 ± 5.24 | 24.00 ± 0.00 | 17.78 ± 5.67 | 32.89 ± 10.31 | 25.86 ± 4.57 |
| GPT-4o* | 25.00 | 25.00 | 22.22 | 20.00 | 56.67 | 38.10 | 39.00 | |
| Claude-3.5* | 16.67 | 25.00 | 33.33 | 25.00 | 66.66 | 66.66 | 47.00 | |
| Claude-3.7 | 45.83 ± 12.50 | 65.00 ± 5.00 | 58.33 ± 2.78 | 46.83 ± 1.50 | 65.00 ± 1.67 | 75.08 ± 3.78 | 60.67 ± 1.23 | |
| smolagents Reflexion | GPT-o3 | 30.56 ± 3.93 | 40.00 ± 0.00 | 44.44 ± 0.00 | 31.44 ± 2.20 | 22.22 ± 8.31 | 39.78 ± 6.14 | 32.33 ± 4.44 |
| Claude-3.7 | 37.50 ± 4.17 | 50.00 ± 0.00 | 83.33 ± 5.56 | 62.67 ± 4.33 | 70.00 ± 10.00 | 64.45 ± 1.39 | 62.81 ± 3.10 | |
| smolagents PDT | GPT-o3 | 8.33 ± 6.80 | 0.00 ± 0.00 | 0.00 ± 0.00 | 3.26 ± 1.30 | 8.62 ± 3.88 | 15.86 ± 5.53 | 7.69 ± 2.69 |
| Claude-3.7 | 16.67 ± 0.00 | 10.00 ± 0.00 | 11.11 ± 0.00 | 11.50 ± 3.50 | 9.24 ± 0.35 | 25.50 ± 5.46 | 14.38 ± 2.15 | |
@@
| System | Metric | Archeology | Astronomy | Biomedical | Environment | Legal | Wildfire | Total |
| DS-Guru few-shot (GPTo3) | Score | 25.00 ± 6.80 | 0.00 ± 0.00 | 3.70 ± 5.24 | 48.33 ± 8.50 | 51.11 ± 3.14 | 48.87 ± 7.90 | 37.84 ± 3.90 |
| Avg. runtime/task (min) | 2.8502 | 1.6584 | 14.4576 | 2.3543 | 3.8403 | 2.3559 | 3.8076 | |
| smolagents DR | Score | 50.00 ± 0.00 | 80.00 ± 0.00 | 55.56 ± 0.00 | 40.83 ± 0.00 | 60.00 ± 0.00 | 67.24 ± 0.00 | 58.12 ± 0.00 |
| Avg. runtime/task (min) | 2.3538 | 5.4944 | 17.9464 | 3.6673 | 3.4209 | 3.2526 | 4.8074 | |
| smolagents Reflexion | Score | 33.33 ± 0.00 | 50.00 ± 0.00 | 66.67 ± 0.00 | 43.33 ± 0.00 | 55.54 ± 0.00 | 64.44 ± 0.00 | 52.81 ± 0.00 |
| Avg. runtime/task (min) | 6.451 | 7.2277 | 43.216 | 7.3128 | 6.5276 | 5.3909 | 9.696 | |
| smolagents PDT | Score | 16.67 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 14.37 ± 0.00 | 13.33 ± 0.00 | 0.00 ± 0.00 | 8.70 ± 0.00 |
| Avg. runtime/task (min) | 7.2708 | 6.3041 | 15.5661 | 6.6759 | 8.031 | 8.5672 | 8.2438 | |
| OpenAI DR | Score | 40.00 | 33.33 | 44.45 | 61.67 | 50.00 | 67.28 | 52.18 |
@@
| SUT | Overall Accuracy | Overall Runtime | Accuracy/ Runtime | Accuracy / 1k In Tokens | Accuracy / 1k Out Tokens |
| GPTo3 - Naive | 05.87 ± 00.71 | 01h 01m 41s ± 06m 26s | 00.17 ± 00.02 | 04.73 ± 00.51 | 02.69 ± 00.31 |
| GPTo3 - One Shot | 16.67 ± 02.93 | 01h 11m 26s ± 08m 13s | 00.42 ± 00.12 | 00.43 ± 00.07 | 09.16 ± 01.63 |
| GPTo3 - Few Shot | 24.98 ± 01.25 | 02h 30m 42s ± 42m 34s | 00.31 ± 00.07 | 00.29 ± 00.05 | 08.58 ± 00.61 |
| GPT4o - Naive | 00.00 ± 00.00 | 38m 42s ± 02m 23s | 00.00 ± 00.00 | 00.00 ± 00.00 | 00.00 ± 00.00 |
| GPT4o - One Shot | 06.08 ± 01.25 | 39m 15s ± 02m 11s | 00.27 ± 00.06 | 00.45 ± 00.09 | 09.81 ± 02.06 |
| GPT4o - Few Shot | 08.67 ± 01.48 | 01h 13m 13s ± 18m 12s | 00.22 ± 00.05 | 00.26 ± 00.03 | 07.68 ± 01.33 |
| Llama3 Instruct - Naive | 00.33 ± 00.46 | 35m 09s ± 03m 53s | 00.01 ± 00.02 | 00.26 ± 00.37 | 00.37 ± 00.52 |
| Llama3 Instruct - One Shot | 03.42 ± 00.37 | 01h 38m 51s ± 11m 06s | 00.06 ± 00.01 | 00.25 ± 00.03 | 05.63 ± 00.61 |
| Llama3 Instruct - Few Shot | 08.40 ± 01.01 | 02h 10m 11s ± 21m 06s | 00.12 ± 00.03 | 00.27 ± 00.04 | 08.98 ± 01.20 |
| DeepseekR1 - Naive | 00.33 ± 00.46 | 01h 36m 29s ± 07m 12s | 00.01 ± 00.01 | 00.26 ± 00.37 | 00.16 ± 00.23 |
| DeepseekR1 - One Shot | 00.00 ± 00.00 | 01h 46m 09s ± 04m 42s | 00.00 ± 00.00 | 00.00 ± 00.00 | 00.00 ± 00.00 |
| DeepseekR1 - Few Shot | 00.33 ± 00.46 | 01h 56m 18s ± 12m 06s | 00.01 ± 00.01 | 00.01 ± 00.01 | 00.15 ± 00.22 |
@@ B.2 Ablation Studies
| SUT | Archeology | Astronomy | Biomedical | Environment | Legal | Wildfire | Overall | Runtime |
@@
| SUT | Archeology | Astronomy | Biomedical | Environment | Legal | Wildfire | Overall | Runtime |
@@
Appendix C smolagents Agentic Baseline Details
In this section, we describe the single-agent and multi-agent baselines systems we evaluated on KramaBench more.
C.1 smolagents DR
For the single-agent baseline, we use the open source deep research implementation by smolagents (69). This agentic framework follows a canonical think action response loop with agentic actions expressed in code. In addition to code, the system is equipped with a text inspector capable of processing different common formats originally released with Microsoft Magentic One (37). While the official implementation also equips the system with a web browser by default, we disabled the internet access to allow for direct comparison with DS-Guru.
C.2 smolagents Reflexion
Our first multi-agent baseline is Reflexion (Shinn et al., 2023). In addition to an Actor agent, Reflexion (Figure 6) introduces (1) An Evaluator agent which provides internal feedback by evaluating the outcome of each action. (2)A Self-reflection agent which provides external feedback with the outcome and the evaluation of the action. Compared to traditional reinforcement learning techniques, feedback in Reflexion are expressed with natural language and stored in agent memory to guide future actions.
C.3 smolagents PDT
Our second multi-agent baseline is based on AutoPrep (Fan et al., 2025), a framework for natural language question answering over tabular data. We augmented AutoPrep with tools for parsing non-tabular data and the hierarchical task decomposition technique similar with DS-Guru to address the complexity of KramaBench tasks. The original AutoPrep pipeline involves three agents: (1) Planner (2) Programmer (3) Executor. With the augmentations we implemented, the system employs two agents respectively playing the roles of (1) Planner and (task) Decomposer (3) Tool Executor. We implemented this approach also using smolagents and illustrate the architecture in Figure 7.
Appendix D Dataset Details
@@
Archeology: the data files consists of chronological, archaeological, faunal, and botanical data supporting the presence of Holocene hunter-gatherers on the Maltese Islands in the Mediterranean from roughly 8000 years ago to 7500 years ago. The files were collected from the publicly available data associated with the papers Groucutt et al. (2021); Scerri et al. (2025).
Astronomy: the data files consist of the OMNI dataset Papitashvili and King (2020a; b) that contains near-Earth solar wind, plasma, and magnetic field data, the Swarm dataset Siemes et al. (2016); European Space Agency (2013) that contains the magnetic field and geomagnetic field data, the SILSO Sunspot Number data Clette and Lefèvre (2015), Space-Track.org Two-Line Element Sets (TLEs) U.S. Space Command (2025), the National Oceanic and Atmospheric Administration (NOAA) Flux Forecast dataset U.S. Air Force and NOAA Space Weather Prediction Center (2025), and NOAA GOES Satellite dataset NOAA Office of Satellite and Product Operations (1994). The combination of these datasets has been used to analyze how activity from the Sun affects Earth’s atmosphere, ocean currents, and weather by the authors of Briden et al. (2023); Parker and Linares (2024).
@@
Environment: the data files consist of beach water quality dataset from Massachusetts Environment Public Health Tracking (EPHT) Massachusetts Department of Public Health (2025a), the Massachusetts Bay beach dataset from Massachusetts Water Resources Authority (MWRA) Massachusetts Water Resources Authority (2025b), and the rainfall dataset from NOAA National Weather Service National Weather Service (2025), from 2002 to 2025. The data has been used in yearly reports Massachusetts Department of Public Health (2025b); Massachusetts Water Resources Authority (2025a) to uncover trends in beach water pollution and the correlation between rainfall and water quality.
@@
Legal: the datasets consists of 136 data files, accessible through the Federal Trade Commission (FTC) portal Federal Trade Commission (2025c) and Wikipedia Wikipedia contributors (2025), including information on merger filings, civil penalty actions, etc. The data is used in visualizations and dashboards that analyze nation-level debt collection and fraud detection, available at Federal Trade Commission (2025b; a).
@@
Wildfire: the datasets consists of NOAA wildfire dataset National Centers for Environmental Information (NCEI) (2025), National Interagency Fire Center (NIFC) Fire Information National Interagency Fire Center (2025), US Environmental Protection Agency (EPA) Air Quality Annual Data U.S. Environmental Protection Agency (2025), US Election 2020 Dataset Fontes (2020), Zillow Home Value Index Dataset Robikscube (2021), US Census 2020 U.S. Census Bureau (2025), and the Large wildfire Incident Status Summary Young et al. (2021) to understand wildfire incident location, cause, and consequences in the US from 2002 to 2016. This data has been used for analysis in the reports published by the NOAA and NIFC NCEI.Monitoring.Info@noaa.gov (2025); Center .
D.1 Cross-domain Accuracy Difference Analysis
In this subsection, we discuss the findings on likely causes of the differences in accuracy between different domains in KramaBench. We obtained these findings by manually analyzing the traces of smolagents-reflexion DR.
1.
Archeology () : In this domain, the system correctly solves questions answerable from a single table. However, errors occur for tasks requiring joining tables found in different files, because it treats multiple files as raw text instead of loading them as tables.
2.
Astronomy () : Astronomy tasks have the lowest average performance. In this domain, a large portion of the required input data is found in proprietary scientific formats (e.g., FORTRAN-style dat files). We observed that the agent struggles whenever it needed to load data from these files, e.g., SP3 orbit files or satellite products.
3.
Biomedical () : When working with biomedical data, the agent is reliable for shallow operations on a single sheet but fails to navigate large, multi-sheet workbooks and join data across sheets. Cross-sheet joins, especially between clinical and phosphoproteomics data, are problematic, and errors arise in correlation statistics due to sign miscalculations.
4.
Environment () : In the environmental domain, the system performs well on the relatively tasks involving clean CSV data, such as filtering, counting, and averaging. Unlike other tasks which struggle from data retrieval or understanding issues, the main issues arise from implementation/arithmetic mistakes, such as incorrect aggregation scopes or rounding errors, which explains the higher score.
5.
Legal (): In these tasks, the agent handles the straightforward pipelines well but struggles with loading data from messy files, i.e., that contain multi-row headers, partial subtotals, and metadata rows. Amongst the common errors that stem from these shortcomings, one example is that sum, means, and aggregations are only partial due to incorrect loading.
6.
Wildfire () : Within wildfire-related tasks, the system faces challenges with geospatial data and temporal/statistical reasoning. It struggles with GeoPackage layers and spatial joins, as well as with rolling-window aggregations for weather. However, text lookups and simple value comparisons work relatively well.
Appendix E Task Details
@@
| Archeology | 12 | 71 | 5 | 2 | 7.5MB | |
| Astronomy | 12 | 68 | 1556 | 8 | 486MB | |
| Biomedical | 9 | 38 | 7 | 2 | 175MB | |
| Environment | 20 | 148 | 37 | 3 | 31MB | |
| Legal | 30 | 188 | 136 | 2 | 1.3MB | |
| Wildfire | 21 | 120 | 23 | 7 | 1GB | |
| Total | 104 | 633 | 1764 | 24 | 1.7GB |
@@
Appendix F Human Baseline Details
In this section, we summarize how we conducted the human baseline and discuss the results and implications.
To contextualize LLM performance on Kramabench, we conducted a human data science study involving nine participants. Each participant was assigned a subset of benchmark tasks and asked to solve them under the same data directory structure, resource constraints, and assumptions provided to our LLM agents. For every assigned task, participants produced:
a complete, reproducible end-to-end solution in a Jupyter notebook,
a detailed log of their active time, broken down into data exploration, pipeline design, coding, and debugging, and
both draft-stage notes and a final clean solution, enabling direct comparison to LLM workflows and error modes.
Participants followed standardized instructions covering repository setup, task management, time tracking, and reproducibility requirements. Across all human-authored solutions, we performed a detailed analysis of the underlying failure causes. The distribution of error types is:
Incorrect pipeline design (): The largest category. These errors occur when, for example, experts mis-specified a join, aggregation rule, grouping key, or filtering logic. This suggests that the most cognitively demanding part of real-world data science is pipeline design, rather than implementation.
Lack of domain knowledge (): Many tasks contain implicit domain assumptions (e.g., definitions of “violation,” mapping categorical labels). Experts often produced internally consistent but mismatched interpretations. This shows that even humans struggle with domain-specific task semantics.
Incorrect inputs (): Tasks often require gathering information across multiple similarly named or structurally similar files, and even humans sometimes use wrong inputs. These errors reflect the challenge of navigating multi-file datasets.
Incorrect answer format (): Some errors are due to having the final outputs in the wrong representation (e.g., units, rounding, formatting), which did not match the one requested by the task.
Library/version issues (): Minor inconsistencies (e.g., pandas handling) that changed intermediate results enough to fail strict correctness checking.
Interpretation and implications.
Multi-file, multi-step pipelines are inherently error-prone—even for trained experts. Humans have difficulty navigating a vast data lake, which we see as an opportunity for LLM-powered systems to quickly search through the lake and identify the target files. Having a reliable retriever could greatly improve accuracy. Ambiguity and assumed domain knowledge are a real factor in real-world data tasks. One possible way for future agentic data-science systems to combat this issue is to ask clarification questions and invite user input. Another approach is to branch out on possible solutions by clearly stating the assumptions. Pipeline design is the bottleneck. Nearly half of all errors () are due to incorrect pipeline logic, highlighting that the core challenge is understanding what transformations to perform, not coding them. Overall, most human errors stemmed from misinterpreting ambiguous tasks, selecting the wrong files, or designing incorrect pipelines—challenges that mirror the dominant failure modes of LLM agents. This confirms that Kramabench captures genuinely difficult, real-world data-to-insight tasks where even trained data scientists struggle with pipeline reasoning, multi-file navigation, and implicit domain assumptions.
Appendix G Evaluation Details